OpenStudy (anonymous):

Solve the following system of equations. 4x – 3y + z = –9 x + 2y + 2z = 5 x – 2y + 3z = –2

OpenStudy (anonymous):

Choices: (1, 1, –10) (1, 2, –7) (–1, 2, 1) (–1, 1, 0)

OpenStudy (anonymous):

Just solved it i think: z = -9 - 4x + 3y x + 2y + 2(-9 - 4x + 3y) = 5 x - 2y + 3(-9 - 4x + 3y) = -2 x + 2y - 18 - 8x + 6y = 5 x - 2y - 27 - 12x + 9y = -2 -7x + 8y = 23 -11x + 7y = 25 4x + y = -2 y = -2 - 4x -7x + 8(-2 - 4x) = 23 -7x - 16 - 32x = 23 -39 x = 39 x = -1 y = -2 - 4x= 2 z = -9 - 4x + 3y = -9 + 4 + 6 = 1 (-1, 2, 1)